### v2

# One-loop Euler-Heisenberg action in Lorentz-violating QED revisited

One-loop Euler-Heisenberg action in Lorentz-violating QED revisited
R. Araújo Universidade Federal de Alagoas, 57072-900, Maceió, Alagoas, Brazil raline.araujo,tmariz@fis.ufal.br    T. Mariz Universidade Federal de Alagoas, 57072-900, Maceió, Alagoas, Brazil raline.araujo,tmariz@fis.ufal.br    J. R. Nascimento Departamento de Física, Universidade Federal da Paraíba,
Caixa Postal 5008, 58051-970, João Pessoa, Paraíba, Brazil
jroberto,petrov@fisica.ufpb.br
   A. Yu. Petrov Departamento de Física, Universidade Federal da Paraíba,
Caixa Postal 5008, 58051-970, João Pessoa, Paraíba, Brazil
jroberto,petrov@fisica.ufpb.br
(10 de outubro de 2025)
Resumo
We discuss applications of the proper-time method in a Lorentz-violating extension of QED characterized by the modification of the mass sector through the addition of the term proportional to the antisymmetric tensor $H_{\mu\nu}$. Unlike other LV extensions of QED, in our case the one-loop Euler-Heisenberg-like action turns out to include only odd degrees of the stress tensor $F_{\mu\nu}$. Our result is shown to be UV finite, and it is confirmed using the Feynman diagrams framework.
I Introduction
An important line of studies of Lorentz-violating (LV) theories consists in obtaining the quantum corrections in these theories. The first example of such a study was performed already in the seminal paper Colladay where the LV Standard Model extension (LV SME) was formulated. Various studies of quantum corrections in LV theories have been performed, with the most important ones are renormalization of LV parameters in LV SME (see e.g. KosPic ) and calculation of finite quantum corrections in LV theories, which is naturally treated as a perturbative generation of LV terms (interesting examples of such calculations are given e.g. in JK ; aether , and a review on obtaining finite corrections to various LV terms is presented in ourLV ).
In this context, one of the interesting problems in studying of effective action in various extensions of QED is certainly the obtaining the one-loop Euler-Heisenberg (EH) effective action EH , which involves all orders in the stress tensor $F_{\mu\nu}$. It has been treated in many contexts (for a review, see Dunne ). Therefore, its computing in LV theories is a rather natural task. The first examples of such calculations are presented in EH1 for certain minimal LV extensions of the spinor QED, and in scalpt for LV scalar QED. However, only a few examples of LV extensions of QED have been studied up to now within the EH context, explicitly, for the spinor LV QED, the LV terms proportional to $b^{\mu}$ and $c^{\mu\nu}$ have been studied in EH1 , while the LV term proportional to $a_{\mu}$ is easily shown to yield a trivial result (these LV tensors are defined in KosPic and listed below in (1)), and it is natural to consider other ones. Within this paper we concentrate on LV modification of the mass sector described through the CPT-even additive term $\frac{1}{2}\bar{\psi}H^{\mu\nu}\sigma_{\mu\nu}\psi$ in the Lagrangian, for which we will find the EH effective action. This is the aim we pursue in this paper.
The structure of the paper is as follows. In Section II, we calculate the EH effective action to the first order in $H_{\mu\nu}$. in Section III, we study the three-point function of the gauge field through the Feynman diagrams framework and explicitly demonstrate that it matches the result that can be read off from the EH effective action. Finally, in Section IV, the discussion of our results is presented.
II Proper-time method
The spinor sector of the minimal LV QED is described by the Lagrangian KosPic
${\cal L}=\bar{\psi}(i\Gamma^{\nu}D_{\nu}-M)\psi,$ (1)
where $D_{\mu}=\partial_{\mu}+ieA_{\mu}$, with
$\Gamma^{\nu}$ $=$ $\gamma^{\nu}+c^{\mu\nu}\gamma_{\mu}+d^{\mu\nu}\gamma_{\mu}\gamma_{5}+e^{\nu}+if^{\nu}\gamma_{5}+{\textstyle\frac{1}{2}}g^{\lambda\mu\nu}\sigma_{\lambda\mu},$
$M$ $=$ $m+a_{\mu}\gamma^{\mu}+b_{\mu}\gamma^{\mu}\gamma_{5}+{\textstyle\frac{1}{2}}H^{\mu\nu}\sigma_{\mu\nu}.$ (2)
Our aim consists in studying the one-loop low-energy effective action in the gauge sector, i.e. the LV EH action. Effectively we must obtain the contribution to the effective action of the lower order in a corresponding LV parameter, but including all orders in fields, from the following functional determinant
$\Gamma^{(1)}$ $=$ $-i{\rm Tr}\ln(i\Gamma^{\nu}D_{\nu}-M).$ (3)
In our earlier studies EH1 , we have obtained the complete one-loop low-energy effective action (LEEA), that is, the EH action, for the lower orders in $b_{\mu}$ and $c_{\mu\nu}$, while the impact of $a^{\mu}$ is trivial in the Abelian case since $a^{\mu}$ is ruled out by the gauge transformation. As argued, and as it is very natural, the lower order in $b_{\mu}$ for the EH action must be the second one (the first order will yield either CFJ term or higher-derivative (HD) terms like Myers-Pospelov or HD CFJ-like terms, see JK ; HDLV , which do not match the EH structure). It is clear also that the EH structure cannot be generated by the first orders in $e^{\mu}$, $f^{\mu}$, $g^{\mu\nu\lambda}$ since they have odd numbers of indices, and by the first order in $d^{\mu\nu}$ for parity reasons, since $d^{\mu\nu}$ is a pseudotensor, so that to yield an observer Lorentz scalar form the quantum correction must include even orders in this parameter.
Therefore, it is natural to study the possible one-loop effective action involving the first order in $H_{\mu\nu}$ which is expected to yield the EH form. While it is clear that the possible EH-like contributions of even orders in $F_{\mu\nu}$, like ${\rm tr}(HF^{2n})$ (e.g. $H_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\phantom{\nu}\alpha}$), vanish by symmetry reasons, and $H^{\mu\nu}F_{\mu\nu}$ is a total derivative, nothing forbids the possibility of terms with higher odd orders in $F_{\mu\nu}$, like ${\rm tr}(HF^{2n+1})$. So, let us study these contributions to the EH action.
We start with the expression (3) which in our case takes the form
[…]
$=$ $-\frac{i}{2}{\rm Tr}\ln[D^{2}+{\textstyle\frac{e}{2}}F_{\mu\nu}\sigma^{\mu\nu}+2H_{\mu\nu}\gamma^{\mu}D^{\nu}+m^{2}+mH_{\mu\nu}\sigma^{\mu\nu}+{\textstyle\frac{1}{4}}(H_{\rho\sigma}\sigma^{\rho\sigma})^{2}],$
where we took into account that $D_{\mu},D_{\nu}]=ieF_{\mu\nu}$. Now, we expand this expression up to the first order in $H^{\mu\nu}$:
$\Gamma^{(1)}$ $=$ $-\frac{i}{2}{\rm Tr}[(2H_{\kappa\lambda}\gamma^{\kappa}D^{\lambda}+mH_{\rho\sigma}\sigma^{\rho\sigma})(D^{2}+{\textstyle\frac{e}{2}}F_{\mu\nu}\sigma^{\mu\nu}+m^{2})^{-1}].$ (6)
To obtain the inverse operator $(D^{2}+{\textstyle\frac{e}{2}}F_{\mu\nu}\sigma^{\mu\nu}+m^{2})^{-1}$, we use the standard proper-time prescription $A^{-1}=-i\int_{0}^{\infty}ds\ e^{isA}$. So, we face a problem of finding $e^{isA}$. Explicitly, our one-loop effective action (6) is
$\Gamma^{(1)}$ $=$ $-\frac{1}{2}{\rm Tr}\int_{0}^{\infty}ds\ e^{ism^{2}}(2H_{\kappa\lambda}\gamma^{\kappa}D^{\lambda}+mH_{\kappa\lambda}\sigma^{\kappa\lambda})\ e^{is(D^{2}+{\textstyle\frac{e}{2}}F_{\mu\nu}\sigma^{\mu\nu})}.$ (7)
As usual, we disregard derivatives of $F_{\mu\nu}$. Thus, to proceed with calculations, we have
$\Gamma^{(1)}$ $=$ $-\frac{1}{2}{\rm tr}\int d^{4}x\int_{0}^{\infty}ds\ e^{ism^{2}}(2H_{\kappa\lambda}\gamma^{\kappa}D^{\lambda}+mH_{\kappa\lambda}\sigma^{\kappa\lambda})\ e^{{\textstyle\frac{ies}{2}}F_{\mu\nu}\sigma^{\mu\nu}}e^{isD^{2}}\delta^{4}(x-x^{\prime})|_{x=x^{\prime}},$ (8)
where we can use the key identity derived earlier in Schwinger:1951nm ; Dittrich:1985yb ; McA :
$e^{isD^{2}}\delta^{4}(x-x^{\prime})|_{x=x^{\prime}}=\langle x|e^{isD^{2}}|x^{\prime}\rangle|_{x=x^{\prime}}=-\frac{i}{16\pi^{2}s^{2}}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$ (9)
Indeed, it is clear that the term proportional to $D^{\lambda}$, in (7), yields a zero contribution to the trace being of odd order in Dirac matrices. Therefore, we are left with
$\Gamma^{(1)}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}H_{\kappa\lambda}\sigma^{\kappa\lambda}\ e^{{\textstyle\frac{ies}{2}}F_{\mu\nu}\sigma^{\mu\nu}}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$ (10)
It remains to expand the exponential in power series in $F_{\mu\nu}$:
$\Gamma^{(1)}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}H_{\kappa\lambda}\sigma^{\kappa\lambda}\sum\limits_{n=0}^{\infty}\frac{({\textstyle\frac{ies}{2}}F_{\mu\nu}\sigma^{\mu\nu})^{2n+1}}{(2n+1)!}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$ (11)
We note that only odd powers of $F_{\mu\nu}$ in the effective action (11) yield nontrivial contributions, since all objects like ${\rm tr}(HF^{2n})$ identically vanish.
It is instructive to calculate the second contribution of (11), namely the term proportional to ${\rm tr}(HF^{3})$, which corresponds to the Lorentz-violating EH action associated with the coefficient $H_{\mu\nu}$. To this end, we must also consider the expansion of the determinant, which reads
${\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right)=1+\frac{e^{2}s^{2}}{12}F_{\mu\nu}F^{\mu\nu}+{\cal O}(F^{4}).$ (12)
Thus, after evaluating the trace, we obtain
$\Gamma_{EH}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}$ (13)
$\times\left[\frac{ie^{3}s^{3}}{24}8H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-\frac{ie^{3}s^{3}}{48}(48H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-64H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa})\right].$
Finally, after performing the $s$-integration, we find
$\Gamma_{EH}=-\frac{e^{3}}{48\pi^{2}m^{3}}\int d^{4}x\ (H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-2H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa}).$ (14)
In the next section we check this result with use of the Feynman diagrams.
Now, in order to study the convergence of the series in (11), we can use the identity
$\{\sigma^{\alpha\beta},\sigma^{\gamma\delta}\}=g^{\alpha\gamma}g^{\beta\delta}-g^{\alpha\delta}g^{\beta\gamma}+i\epsilon^{\alpha\beta\gamma\delta}\gamma_{5},$ (15)
which implies
$(F_{\mu\nu}\sigma^{\mu\nu})^{2}=F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu}\gamma_{5}.$ (16)
Hence, by also taking into account that the term with $n=0$ is trivial and can be disregarded, we can rewrite (11) as
$\Gamma^{(1)}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}H_{\kappa\lambda}\sigma^{\kappa\lambda}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}(F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu}\gamma_{5})^{n}\ F_{\rho\tau}\sigma^{\rho\tau}$ (17)
$\times\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
Then, using the general formula
$(a+b\gamma_{5})^{k}=\frac{(a+b)^{k}+(a-b)^{k}}{2}+\gamma_{5}\frac{(a+b)^{k}-(a-b)^{k}}{2},$ (18)
valid for any positive integer $k$, and taking $a=F_{\mu\nu}F^{\mu\nu}$, and $b=iF_{\mu\nu}\tilde{F}^{\mu\nu}$, we obtain
$\Gamma^{(1)}$ $=$ $\frac{im}{64\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}H_{\kappa\lambda}\sigma^{\kappa\lambda}F_{\rho\tau}\sigma^{\rho\tau}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}$ (19)
$\times[(F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}+(F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}$
$+\gamma_{5}(F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}-\gamma_{5}(F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}]$
$\times\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
It now remains to calculate the traces:
${\rm tr}(\sigma^{\mu\nu}\sigma^{\rho\sigma})=4(g^{\mu\rho}g^{\nu\sigma}-g^{\mu\sigma}g^{\nu\rho}),$
${\rm tr}(\sigma^{\mu\nu}\sigma^{\rho\sigma}\gamma_{5})=-4i\epsilon^{\mu\nu\rho\sigma}.$ (20)
So, we obtain
$\Gamma^{(1)}$ $=$ $\frac{im}{16\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}$ (21)
$\times\{H_{\kappa\lambda}F^{\kappa\lambda}[(F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}+(F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}]$
$-i\epsilon^{\kappa\lambda\rho\tau}H_{\kappa\lambda}F_{\rho\tau}[(F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}-(F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}]\}$
$\times\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
This sum can be evaluated. We note that the result is finite (indeed, the sum begins with $s^{3}$) as it must be by dimensional reasons.
To perform the sum, we introduce the notation: $\Sigma=F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu}$, $\bar{\Sigma}=F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu}$. So, we can write
$\Gamma^{(1)}$ $=$ $\frac{im}{16\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}[H_{\kappa\lambda}F^{\kappa\lambda}(\Sigma^{n}+\bar{\Sigma}^{n})-i\epsilon^{\kappa\lambda\rho\tau}H_{\kappa\lambda}F_{\rho\tau}(\Sigma^{n}-\bar{\Sigma}^{n})]$ (22)
$\times\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
Using the Taylor series $\sum\limits_{n=1}^{\infty}\frac{A^{n}}{(2n+1)!}=\frac{\sinh\sqrt{A}-\sqrt{A}}{\sqrt{A}}$, we finally obtain
$\Gamma^{(1)}$ $=$ $-\frac{em}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s}e^{ism^{2}}\left[H_{\kappa\lambda}F^{\kappa\lambda}\left(\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\Sigma})-{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}{{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}+\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}})-{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}{{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}\right)\right$. (23)
$\left.+i\epsilon^{\kappa\lambda\rho\tau}H_{\kappa\lambda}F_{\rho\tau}\left(\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\Sigma})-{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}{{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}-\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}})-{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}{{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}\right)\right]\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
This is our final result. As we already noted, it is real, UV finite and begins with the cubic order in the stress tensor $F_{\mu\nu}$, which is nontrivial since earlier, in EH1 ; scalpt , only even orders in the stress tensor were shown to yield nontrivial contributions to the effective action.
III Three-point function
In this section, we analyze the three-point function to calculate the one-loop EH effective action of the parameter $H_{\mu\nu}$ using the Feynman diagram approach. Our aim is to explicitly demonstrate that the result obtained reproduces the same structure that can be derived directly from the EH effective action (11).
For this, given the Lagrangian
[…]
we must expand the propagator
[…]
$\Gamma^{(1,3)}$ $=$ $\frac{1}{3}\int d^{4}x\int d^{4}k_{1}d^{4}k_{2}d^{4}k_{3}\ e^{-i(k_{1}+k_{2}+k_{3})\cdot x}\frac{1}{4}G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})A_{\mu_{1}}(k_{1})A_{\mu_{2}}(k_{2})A_{\mu_{3}}(k_{3}),$ (26)
with
$G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=2T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+2T_{H}^{\mu_{2}\mu_{1}\mu_{3}}(k_{2},k_{1},k_{3})$ (27)
where $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ is subdivided into three graphs as follows:
$T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (28)
with
$iT_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-(-{\textstyle\frac{i}{2}})(-ie)^{3}\int\frac{d^{4}p}{(2\pi)^{4}}\mathrm{tr}\ iS(p)H_{\alpha\beta}\sigma^{\alpha\beta}iS(p)\gamma^{\mu_{1}}iS(p_{1})\gamma^{\mu_{2}}iS(p_{12})\gamma^{\mu_{3}},\$(29a)
$iT_{H2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-(-{\textstyle\frac{i}{2}})(-ie)^{3}\int\frac{d^{4}p}{(2\pi)^{4}}\mathrm{tr}\ iS(p)\gamma^{\mu_{1}}iS(p_{1})H_{\alpha\beta}\sigma^{\alpha\beta}iS(p_{1})\gamma^{\mu_{2}}iS(p_{12})\gamma^{\mu_{3}},\$(29b)
$iT_{H3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-(-{\textstyle\frac{i}{2}})(-ie)^{3}\int\frac{d^{4}p}{(2\pi)^{4}}\mathrm{tr}\ iS(p)\gamma^{\mu_{1}}iS(p_{1})\gamma^{\mu_{2}}iS(p_{12})H_{\alpha\beta}\sigma^{\alpha\beta}iS(p_{12})\gamma^{\mu_{3}},\$(29c)
where $p_{1}^{\mu}=p^{\mu}-k_{1}^{\mu}$ and $p_{12}^{\mu}=p^{\mu}-k_{1}^{\mu}-k_{2}^{\mu}$. It is easy to see that $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ can be also obtained when we perform the ciclic interchanges:
$T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $T_{H1}^{\mu_{2}\mu_{3}\mu_{1}}(k_{2},k_{3},k_{1}),$ (30a)
$T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $T_{H1}^{\mu_{3}\mu_{1}\mu_{2}}(k_{3},k_{1},k_{2}).$ (30b)
Therefore, we must focus our attention on the graph $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$, in which, by considering first the Feynman parameterization, we obtain
$T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\int_{0}^{1}dx_{1}\int_{0}^{1-x_{1}}dx_{2}\int\frac{d^{4}p}{(2\pi)^{4}}\frac{3ixe^{3}}{(p^{2}-M^{2})^{4}}$ (31)
[…]
where
$M^{2}=m^{2}-x_{1}(1-x_{1})k_{1}^{2}-x_{12}(1-x_{12})k_{2}^{2}-2x_{1}(1-x_{12})k_{1}\cdot k_{2}$ (32)
and
$q^{\mu}$ $=$ $p^{\mu}+(1-x_{1})k_{1}^{\mu}+(1-x_{12})k_{2}^{\mu},$ (33)
$q_{1}^{\mu}$ $=$ $p_{1}^{\mu}+(1-x_{1})k_{1}^{\mu}+(1-x_{12})k_{2}^{\mu},$ (34)
with $x_{12}=x_{1}+x_{2}$, and so on. Then, after we calculate the trace over Dirac matrices and the corresponding integrals, up to order $1/m^{3}$, we arrive at
$T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=\sum_{i=1}^{4}T_{H1ggi}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}T_{H1gki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}T_{H1kki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (35)
where
$T_{H1gg1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{48\pi^{2}m}(g^{\mu_{1}\mu_{2}}((k_{2}^{\beta}-2k_{1}^{\beta})g^{\alpha\mu_{3}}+(2k_{1}^{\alpha}-k_{2}^{\alpha})g^{\beta\mu_{3}})+g^{\mu_{1}\mu_{3}}(k_{2}^{\alpha}g^{\beta\mu_{2}}-k_{2}^{\beta}g^{\alpha\mu_{2}})$ (36a)
$+g^{\mu_{2}\mu_{3}}((2k_{1}^{\beta}+3k_{2}^{\beta})g^{\alpha\mu_{1}}-(2k_{1}^{\alpha}+3k_{2}^{\alpha})g^{\beta\mu_{1}})),$
$T_{H1gg2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}g^{\mu_{1}\mu_{2}}(-2k_{1}^{\alpha}(2k_{1}^{2}+4k_{2}^{2}+7k_{1}\cdot k_{2})g^{\beta\mu_{3}}+k_{2}^{\alpha}(11k_{1}^{2}+2k_{2}^{2}+6k_{1}\cdot k_{2})g^{\beta\mu_{3}}$ (36b)
$+g^{\alpha\mu_{3}}(2k_{1}^{\beta}(2k_{1}^{2}+4k_{2}^{2}+7k_{1}\cdot k_{2})-k_{2}^{\beta}(11k_{1}^{2}+2k_{2}^{2}+6k_{1}\cdot k_{2}))),$
$T_{H1gg3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}g^{\mu_{1}\mu_{3}}(5k_{1}^{\alpha}(k_{2}^{2}+2k_{1}\cdot k_{2})g^{\beta\mu_{2}}-k_{2}^{\alpha}(11k_{1}^{2}+2k_{2}^{2}+6k_{1}\cdot k_{2})g^{\beta\mu_{2}}$ (36c)
$+g^{\alpha\mu_{2}}(k_{2}^{\beta}(11k_{1}^{2}+2k_{2}^{2}+6k_{1}\cdot k_{2})-5k_{1}^{\beta}(k_{2}^{2}+2k_{1}\cdot k_{2}))),$
$T_{H1gg4}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}g^{\mu_{2}\mu_{3}}(2k_{1}^{\alpha}(-2k_{1}^{2}+k_{2}^{2}+3k_{1}\cdot k_{2})g^{\beta\mu_{1}}-5k_{2}^{\alpha}(3k_{1}^{2}+k_{2}^{2}+2k_{1}\cdot k_{2})g^{\beta\mu_{1}}$ (36d)
$+g^{\alpha\mu_{1}}(k_{1}^{\beta}(4k_{1}^{2}-2k_{2}^{2}-6k_{1}\cdot k_{2})+5k_{2}^{\beta}(3k_{1}^{2}+k_{2}^{2}+2k_{1}\cdot k_{2}))),$
$T_{H1gk1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{48\pi^{2}m}(g^{\alpha\mu_{1}}((k_{2}^{\mu_{3}}-2k_{1}^{\mu_{3}})g^{\beta\mu_{2}}+3(2k_{1}^{\mu_{2}}+k_{2}^{\mu_{2}})g^{\beta\mu_{3}})-g^{\alpha\mu_{2}}((k_{2}^{\mu_{3}}-2k_{1}^{\mu_{3}})g^{\beta\mu_{1}}$ (36e)
$+(2k_{1}^{\mu_{1}}+3k_{2}^{\mu_{1}})g^{\beta\mu_{3}})+g^{\alpha\mu_{3}}((2k_{1}^{\mu_{1}}+3k_{2}^{\mu_{1}})g^{\beta\mu_{2}}-3(2k_{1}^{\mu_{2}}+k_{2}^{\mu_{2}})g^{\beta\mu_{1}})),$
$T_{H1gk2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{96\pi^{2}m^{3}}(g^{\mu_{1}\mu_{2}}(2k_{1}^{\mu_{3}}+k_{2}^{\mu_{3}})-g^{\mu_{1}\mu_{3}}(2k_{1}^{\mu_{2}}+k_{2}^{\mu_{2}})+g^{\mu_{2}\mu_{3}}(2k_{1}^{\mu_{1}}+k_{2}^{\mu_{1}}))$ (36f)
$\times(k_{2}^{\alpha}k_{1}^{\beta}-k_{1}^{\alpha}k_{2}^{\beta}),$
$T_{H1gk3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}(g^{\alpha\mu_{1}}(g^{\beta\mu_{2}}(k_{1}^{\mu_{3}}(4k_{1}^{2}+9k_{2}^{2}+8k_{1}\cdot k_{2})-k_{2}^{\mu_{3}}(11k_{1}^{2}+k_{2}^{2}+7k_{1}\cdot k_{2}))$ (36g)
$-5g^{\beta\mu_{3}}(k_{1}^{\mu_{2}}(3k_{2}^{2}+4(k_{1}^{2}+k_{1}\cdot k_{2}))+k_{2}^{\mu_{2}}(2k_{1}^{2}+k_{2}^{2}+k_{1}\cdot k_{2})))$
$+g^{\alpha\mu_{2}}g^{\beta\mu_{3}}((4k_{1}^{2}+5k_{2}^{2})k_{1}^{\mu_{1}}+5k_{2}^{\mu_{1}}(2k_{2}^{2}+3(k_{1}^{2}+k_{1}\cdot k_{2})))),$
$T_{H1gk4}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}(g^{\alpha\mu_{2}}g^{\beta\mu_{1}}(k_{1}^{\mu_{3}}(4k_{1}^{2}+9k_{2}^{2}+8k_{1}\cdot k_{2})-k_{2}^{\mu_{3}}(11k_{1}^{2}+k_{2}^{2}+7k_{1}\cdot k_{2}))$ (36h)
$+g^{\alpha\mu_{3}}(g^{\beta\mu_{2}}((4k_{1}^{2}+5k_{2}^{2})k_{1}^{\mu_{1}}+5k_{2}^{\mu_{1}}(2k_{2}^{2}+3(k_{1}^{2}+k_{1}\cdot k_{2})))$
$-5g^{\beta\mu_{1}}(k_{1}^{\mu_{2}}(3k_{2}^{2}+4(k_{1}^{2}+k_{1}\cdot k_{2}))+k_{2}^{\mu_{2}}(2k_{1}^{2}+k_{2}^{2}+k_{1}\cdot k_{2})))),$
$T_{H1kk1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{60\pi^{2}m^{3}}(k_{1}^{\mu_{2}}k_{1}^{\mu_{1}}((k_{1}^{\beta}+k_{2}^{\beta})g^{\alpha\mu_{3}}-(k_{1}^{\alpha}+k_{2}^{\alpha})g^{\beta\mu_{3}})+k_{1}^{\mu_{3}}(k_{1}^{\mu_{1}}(k_{2}^{\alpha}g^{\beta\mu_{2}}-k_{2}^{\beta}g^{\alpha\mu_{2}})$ (36i)
$+k_{1}^{\mu_{2}}(k_{1}^{\alpha}g^{\beta\mu_{1}}-k_{1}^{\beta}g^{\alpha\mu_{1}}))),$
$T_{H1kk2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}(k_{2}^{\mu_{2}}k_{1}^{\mu_{1}}((k_{2}^{\beta}-4k_{1}^{\beta})g^{\alpha\mu_{3}}+(4k_{1}^{\alpha}-k_{2}^{\alpha})g^{\beta\mu_{3}})+k_{2}^{\mu_{3}}(k_{1}^{\mu_{1}}((10k_{1}^{\alpha}+k_{2}^{\alpha})g^{\beta\mu_{2}}$ (36j)
$-(10k_{1}^{\beta}+k_{2}^{\beta})g^{\alpha\mu_{2}})+k_{1}^{\mu_{2}}((14k_{1}^{\beta}+5k_{2}^{\beta})g^{\alpha\mu_{1}}-(14k_{1}^{\alpha}+5k_{2}^{\alpha})g^{\beta\mu_{1}}))),$
$T_{H1kk3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{480\pi^{2}m^{3}}(k_{1}^{\mu_{2}}k_{2}^{\mu_{1}}((6k_{1}^{\alpha}+k_{2}^{\alpha})g^{\beta\mu_{3}}-(6k_{1}^{\beta}+k_{2}^{\beta})g^{\alpha\mu_{3}})+k_{1}^{\mu_{3}}(k_{2}^{\mu_{1}}((10k_{1}^{\beta}+k_{2}^{\beta})g^{\alpha\mu_{2}}$ (36k)
$-(10k_{1}^{\alpha}+k_{2}^{\alpha})g^{\beta\mu_{2}})+k_{2}^{\mu_{2}}((5k_{2}^{\beta}-4k_{1}^{\beta})g^{\alpha\mu_{1}}+(4k_{1}^{\alpha}-5k_{2}^{\alpha})g^{\beta\mu_{1}}))),$
$T_{H1kk4}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{240\pi^{2}m^{3}}(2k_{2}^{\mu_{1}}k_{2}^{\mu_{2}}((k_{1}^{\alpha}+k_{2}^{\alpha})g^{\beta\mu_{3}}-(k_{1}^{\beta}+k_{2}^{\beta})g^{\alpha\mu_{3}})-k_{2}^{\mu_{3}}(k_{2}^{\mu_{1}}((5k_{1}^{\alpha}+2k_{2}^{\alpha})g^{\beta\mu_{2}}$ (36l)
$-(5k_{1}^{\beta}+2k_{2}^{\beta})g^{\alpha\mu_{2}})+3k_{2}^{\mu_{2}}(k_{1}^{\beta}g^{\alpha\mu_{1}}-k_{1}^{\alpha}g^{\beta\mu_{1}}))).$
Let us now briefly discuss the question of photon double splitting in the collinear limit. If the incident on-shell photon has energy $E_{1}$ and momentum $\vec{k}_{1}$, then energy–momentum conservation requires that all momenta $\vec{k}_{i}$ must be aligned. Thus, the initial and final photons propagate collinearly, with orthogonal four-momenta $k_{i}^{\mu}k_{j\mu}=0$ (for more details, see Kostelecky:2002ue ). From the transversality condition, one finds $\epsilon_{i\mu}k_{j}^{\mu}=0$ (or $A_{\mu}(k_{i})k_{j}^{\mu}=0$). Hence, the contribution (36a) of order $\frac{1}{m}$ can yield a nonzero amplitude, while all other terms of (36) of order $\frac{1}{m^{3}}$ vanish in this limit. However, when we consider the other contributions coming from $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ through the permutations (30), all the contributions of order $\frac{1}{m}$ cancel each other, and thus photon double splitting does not occur. This fact was only mentioned in Kostelecky:2002ue , but not explicitly shown.
For the photon triple splitting initially discussed in Kostelecky:2002ue , and later in FuMa , the contributions of order $\frac{1}{m^{2}}$ do not cancel each other when all the insertions of the coefficient $c_{\mu\nu}$ in the propagator are considered. Thus, after the collinear limit is taken into account, a nonzero amplitude is obtained, in contrast to what we observe here for the coefficient $H_{\mu\nu}$.
The next step is to analyze the generation of the Lorentz-violating EH action, for the first order correction. For this, we must calculate $G_{H}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq. (27)), by initially calculating $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq. (28)), when we interchange the uncontracted indices as well as the momentum indices for obtaining $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq. (30)) from $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq. (36)). Proceeding in this way, we obtain
$G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=\sum_{i=1}^{3}G_{Hggi}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{3}G_{Hgki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}G_{Hkki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (37)
where
$G_{Hgg1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}g^{\mu_{1}\mu_{2}}(k_{1}^{\alpha}(k_{2}^{2}-k_{1}\cdot k_{2})g^{\beta\mu_{3}}+k_{2}^{\alpha}(k_{1}^{2}-k_{1}\cdot k_{2})g^{\beta\mu_{3}}+g^{\alpha\mu_{3}}$ (38a)
$\times(k_{1}^{\beta}(k_{1}\cdot k_{2}-k_{2}^{2})+k_{2}^{\beta}(k_{1}\cdot k_{2}-k_{1}^{2}))),$
$G_{Hgg2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}g^{\mu_{1}\mu_{3}}(-(k_{1}^{\alpha}(k_{2}^{2}+2k_{1}\cdot k_{2})g^{\beta\mu_{2}})+k_{2}^{\alpha}(2k_{1}^{2}+k_{1}\cdot k_{2})g^{\beta\mu_{2}}$ (38b)
$+g^{\alpha\mu_{2}}(k_{1}^{\beta}(k_{2}^{2}+2k_{1}\cdot k_{2})-k_{2}^{\beta}(2k_{1}^{2}+k_{1}\cdot k_{2}))),$
$G_{Hgg3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}g^{\mu_{2}\mu_{3}}(-(k_{1}^{\alpha}(2k_{2}^{2}+k_{1}\cdot k_{2})g^{\beta\mu_{1}})+k_{2}^{\alpha}(k_{1}^{2}+2k_{1}\cdot k_{2})g^{\beta\mu_{1}}$ (38c)
$+g^{\alpha\mu_{1}}(k_{1}^{\beta}(2k_{2}^{2}+k_{1}\cdot k_{2})-k_{2}^{\beta}(k_{1}^{2}+2k_{1}\cdot k_{2}))),$
$G_{Hgk1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}(g^{\mu_{1}\mu_{2}}(k_{1}^{\mu_{3}}-k_{2}^{\mu_{3}})-g^{\mu_{1}\mu_{3}}(2k_{1}^{\mu_{2}}+k_{2}^{\mu_{2}})+g^{\mu_{2}\mu_{3}}(k_{1}^{\mu_{1}}+2k_{2}^{\mu_{1}}))$ (38d)
$\times(k_{2}^{\alpha}k_{1}^{\beta}-k_{1}^{\alpha}k_{2}^{\beta}),$
$G_{Hgk2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}(g^{\alpha\mu_{1}}(g^{\beta\mu_{2}}(k_{1}^{\mu_{3}}(k_{2}^{2}+k_{1}\cdot k_{2})-k_{2}^{\mu_{3}}(k_{1}^{2}+k_{1}\cdot k_{2}))+g^{\beta\mu_{3}}$ (38e)
$\times(k_{2}^{2}k_{1}^{\mu_{2}}-k_{2}^{\mu_{2}}k_{1}\cdot k_{2}))+g^{\alpha\mu_{2}}g^{\beta\mu_{3}}(k_{1}^{2}k_{2}^{\mu_{1}}-k_{1}^{\mu_{1}}k_{1}\cdot k_{2})),$
$G_{Hgk3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}(g^{\alpha\mu_{2}}g^{\beta\mu_{1}}(k_{1}^{\mu_{3}}(k_{2}^{2}+k_{1}\cdot k_{2})-k_{2}^{\mu_{3}}(k_{1}^{2}+k_{1}\cdot k_{2}))+g^{\alpha\mu_{3}}$ (38f)
$\times(g^{\beta\mu_{1}}(k_{2}^{2}k_{1}^{\mu_{2}}-k_{2}^{\mu_{2}}k_{1}\cdot k_{2})+g^{\beta\mu_{2}}(k_{1}^{2}k_{2}^{\mu_{1}}-k_{1}^{\mu_{1}}k_{1}\cdot k_{2}))),$
$G_{Hkk1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}(k_{1}^{\mu_{2}}k_{1}^{\mu_{1}}(k_{2}^{\beta}g^{\alpha\mu_{3}}-k_{2}^{\alpha}g^{\beta\mu_{3}})+k_{1}^{\mu_{3}}(2k_{1}^{\mu_{1}}(k_{2}^{\alpha}g^{\beta\mu_{2}}-k_{2}^{\beta}g^{\alpha\mu_{2}})$ (38g)
$+k_{1}^{\mu_{2}}(k_{2}^{\beta}g^{\alpha\mu_{1}}-k_{2}^{\alpha}g^{\beta\mu_{1}}))),$
$G_{Hkk2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}k_{2}^{\mu_{3}}(k_{1}^{\mu_{1}}(k_{1}^{\beta}g^{\alpha\mu_{2}}-k_{1}^{\alpha}g^{\beta\mu_{2}})+k_{1}^{\mu_{2}}((k_{2}^{\beta}-k_{1}^{\beta})g^{\alpha\mu_{1}}$ (38h)
$+(k_{1}^{\alpha}-k_{2}^{\alpha})g^{\beta\mu_{1}})),$
$G_{Hkk3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}(k_{1}^{\mu_{2}}k_{2}^{\mu_{1}}((k_{1}^{\alpha}+k_{2}^{\alpha})g^{\beta\mu_{3}}-(k_{1}^{\beta}+k_{2}^{\beta})g^{\alpha\mu_{3}})+k_{1}^{\mu_{3}}(k_{2}^{\mu_{1}}((k_{1}^{\beta}-k_{2}^{\beta})g^{\alpha\mu_{2}}$ (38i)
$+(k_{2}^{\alpha}-k_{1}^{\alpha})g^{\beta\mu_{2}})+k_{2}^{\mu_{2}}(k_{2}^{\beta}g^{\alpha\mu_{1}}-k_{2}^{\alpha}g^{\beta\mu_{1}}))),$
$G_{Hkk4}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}(k_{2}^{\mu_{2}}k_{2}^{\mu_{1}}(k_{1}^{\alpha}g^{\beta\mu_{3}}-k_{1}^{\beta}g^{\alpha\mu_{3}})+k_{2}^{\mu_{3}}(k_{2}^{\mu_{1}}(k_{1}^{\alpha}g^{\beta\mu_{2}}-k_{1}^{\beta}g^{\alpha\mu_{2}})$ (38j)
$+2k_{2}^{\mu_{2}}(k_{1}^{\beta}g^{\alpha\mu_{1}}-k_{1}^{\alpha}g^{\beta\mu_{1}}))).$
We observe that the contributions of order $\frac{1}{m}$ vanish, as expected.
Thus, by considering these results (37), the effective action (26) takes the form
$\Gamma_{EH}$ $=$ $\int d^{4}x\int d^{4}k_{1}d^{4}k_{2}d^{4}k_{3}\ e^{-i(k_{1}+k_{2}+k_{3})\cdot x}G_{H}(k_{1},k_{2},k_{3}),$ (39)
where
$G_{H}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}}{144\pi^{2}m^{3}}(H_{\kappa\lambda}F^{\kappa\lambda}(k_{1})F_{\mu\nu}(k_{2})F^{\mu\nu}(k_{3})+H_{\kappa\lambda}F^{\kappa\lambda}(k_{2})F_{\mu\nu}(k_{1})F^{\mu\nu}(k_{3})$ (40)
$+H_{\kappa\lambda}F^{\kappa\lambda}(k_{3})F_{\mu\nu}(k_{1})F^{\mu\nu}(k_{2})-2H_{\kappa\lambda}F^{\lambda\mu}(k_{1})F_{\mu\nu}(k_{2})F^{\nu\kappa}(k_{3})$
$-2F_{\kappa\lambda}(k_{1})H^{\lambda\mu}F_{\mu\nu}(k_{2})F^{\nu\kappa}(k_{3})-2F_{\kappa\lambda}(k_{1})F^{\lambda\mu}(k_{2})H_{\mu\nu}F^{\nu\kappa}(k_{3})),$
with $F^{\mu\nu}(k_{1})=k_{1}^{\mu}A^{\nu}(k_{1})-k_{1}^{\nu}A^{\mu}(k_{1})$, and so on. Then, inverting the Fourier transform in Eq. (39), the Lorentz-violating Euler-Heisenberg action becomes
$\Gamma_{EH}=-\frac{e^{3}}{48\pi^{2}m^{3}}\int d^{4}x\ (H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-2H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa}),$ (41)
where now $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. This result exactly coincides with the expression (14) obtained above with use of the proper time methods which confirms the validity of our calculations.
IV Summary
We calculated the one-loop EH effective action for the LV spinor QED characterized by the LV parameter $H_{\mu\nu}$ of dimension one, which ensures finiteness of our result. Its unusual feature consists in the fact that, unlike the standard EH effective action (including many cases with the presence of LV terms, see EH1 ; scalpt ), our one-loop expression involves only odd orders in $F_{\mu\nu}$. Another important observation regarding our result is that it involves the first order in the LV parameter while in many other studies of the EH effective action the result is of the second order in corresponding LV parameters (see e.g. EH1 ; scalpt ). Therefore, our result may be the dominant contribution due to the well-known smallness of LV parameters datatables . We explicitly demonstrated that the third-order result obtained through the EH framework coincides with that calculated through the Feynman diagrams approach. Effectively, we succeeded in obtaining results analogous to those found in FuMa ; Kostelecky:2002ue for our theory. We plan to study phenomenological impacts of our results in further studies.
Another interesting problem consists in obtaining the EH effective action depending on other minimal LV parameters presented in the model (1), namely, $d^{\mu\nu},e^{\mu},f^{\mu},g^{\mu\nu\lambda}$. We note that these calculations will be more complicated than those performed here and in EH1 , since in the presence of these parameters, calculations of more involved traces will be necessary. We plan to perform these studies in our forthcoming papers.
Acknowledgments. The work of T. M. has been partially supported by the CNPq project No. 309360/2025-0 and FAPEAL project No. E:60030.0000002341/2022. The work of A. Yu. P. has been partially supported by the CNPq project No. 303777/2023-0.
Referências
  • (1) D. Colladay and V. A. Kostelecky, Phys. Rev. D 58 (1998), 116002 [arXiv:hep-ph/9809521 [hep-ph]].
  • (2) V. A. Kostelecky, C. D. Lane and A. G. M. Pickering, Phys. Rev. D 65 (2002), 056006 [arXiv:hep-th/0111123 [hep-th]].
  • (3) R. Jackiw and V. A. Kostelecky, Phys. Rev. Lett. 82 (1999), 3572-3575 [arXiv:hep-ph/9901358 [hep-ph]].
  • (4) M. Gomes, J. R. Nascimento, A. Y. Petrov and A. J. da Silva, Phys. Rev. D 81 (2010), 045018 [arXiv:0911.3548 [hep-th]].
  • (5) A. F. Ferrari, J. R. Nascimento and A. Y. Petrov, Eur. Phys. J. C 80 (2020) no.5, 459 [arXiv:1812.01702 [hep-th]].
  • (6) W. Heisenberg and H. Euler, Z. Phys. 98 (1936) no.11-12, 714-732 [arXiv:physics/0605038 [physics]].
  • (7) G. V. Dunne, [arXiv:hep-th/0406216 [hep-th]].
  • (8) A. F. Ferrari, J. Furtado, J. F. Assunção, T. Mariz and A. Y. Petrov, EPL 136 (2021) no.2, 21002 [arXiv:2109.11901 [hep-th]].
  • (9) L. C. T. Brito, J. C. C. Felipe, A. C. Lehum, J. R. Nascimento and A. Y. Petrov, Int. J. Mod. Phys. A 40 (2025) no.20, 2550067 [arXiv:2405.20855 [hep-th]].
  • (10) T. Mariz, J. R. Nascimento and A. Y. Petrov, Phys. Rev. D 85 (2012), 125003 [arXiv:1111.0198 [hep-th]].
  • (11) J. S. Schwinger, Phys. Rev. 82 (1951), 664-679.
  • (12) W. Dittrich and M. Reuter, Lect. Notes Phys. 220 (1985), 1-244.
  • (13) I. N. McArthur and T. D. Gargett, Nucl. Phys. B 497 (1997), 525-540. [arXiv:hep-th/9705200 [hep-th]].
  • (14) V. A. Kostelecky and N. Russell, Rev. Mod. Phys. 83 (2011), 11-31 [arXiv:0801.0287 [hep-ph]].
  • (15) V. A. Kostelecky and A. G. M. Pickering, Phys. Rev. Lett. 91 (2003), 031801 [arXiv:hep-ph/0212382 [hep-ph]].
  • (16) J. Furtado and T. Mariz, Phys. Rev. D 89 (2014) no.2, 025021 [arXiv:1401.0492 [hep-ph]].

### v3

@@
 jroberto,petrov@fisica.ufpb.br
-(10 de outubro de 2025)
+(23 de janeiro de 2026)
 Resumo
-We discuss applications of the proper-time method in a Lorentz-violating extension of QED characterized by the modification of the mass sector through the addition of the term proportional to the antisymmetric tensor $H_{\mu\nu}$.
-Unlike other LV extensions of QED, in our case the one-loop Euler-Heisenberg-like action turns out to include only odd degrees of the stress tensor $F_{\mu\nu}$.
+We discuss applications of the proper-time method in a Lorentz-violating extension of QED characterized by the addition of the term proportional to the antisymmetric tensor $H_{\mu\nu}$.
+Unlike other LV extensions of QED, in our case, the one-loop Euler-Heisenberg-like action turns out to include only odd powers of the stress tensor $F_{\mu\nu}$.
 Our result is shown to be UV finite, and it is confirmed using the Feynman diagrams framework.
 I Introduction
-An important line of studies of Lorentz-violating (LV) theories consists in obtaining the quantum corrections in these theories.
+An important line of studies of Lorentz-violating (LV) theories consists of obtaining the quantum corrections in these theories.
 The first example of such a study was performed already in the seminal paper Colladay where the LV Standard Model extension (LV SME) was formulated.
-Various studies of quantum corrections in LV theories have been performed, with the most important ones are renormalization of LV parameters in LV SME (see e.g.
+Various studies of quantum corrections in LV theories have been performed, where the most important ones are renormalization of LV parameters in LV SME (see e.g.
 KosPic ) and calculation of finite quantum corrections in LV theories, which is naturally treated as a perturbative generation of LV terms (interesting examples of such calculations are given e.g.
 in JK ; aether , and a review on obtaining finite corrections to various LV terms is presented in ourLV ).
-In this context, one of the interesting problems in studying of effective action in various extensions of QED is certainly the obtaining the one-loop Euler-Heisenberg (EH) effective action EH , which involves all orders in the stress tensor $F_{\mu\nu}$.
+In this context, one of the interesting problems in studying the effective action in various extensions of QED is certainly to obtain the one-loop Euler-Heisenberg (EH) effective action EH , which involves all orders in the stress tensor $F_{\mu\nu}$.
 It has been treated in many contexts (for a review, see Dunne ).
-Therefore, its computing in LV theories is a rather natural task.
+Therefore, computing it in LV theories is a rather natural task.
 The first examples of such calculations are presented in EH1 for certain minimal LV extensions of the spinor QED, and in scalpt for LV scalar QED.
-However, only a few examples of LV extensions of QED have been studied up to now within the EH context, explicitly, for the spinor LV QED, the LV terms proportional to $b^{\mu}$ and $c^{\mu\nu}$ have been studied in EH1 , while the LV term proportional to $a_{\mu}$ is easily shown to yield a trivial result (these LV tensors are defined in KosPic and listed below in (1)), and it is natural to consider other ones.
-Within this paper we concentrate on LV modification of the mass sector described through the CPT-even additive term $\frac{1}{2}\bar{\psi}H^{\mu\nu}\sigma_{\mu\nu}\psi$ in the Lagrangian, for which we will find the EH effective action.
+However, only a few examples of LV extensions of QED have been studied up to now within the EH context.
+Explicitly, for the spinor LV QED, the LV terms proportional to $b^{\mu}$ and $c^{\mu\nu}$ have been studied in EH1 .
+At the same time, the LV term proportional to $a_{\mu}$ is easily shown to yield a trivial result (these LV tensors are defined in KosPic and listed below in (1)).
+So, it is natural to consider others.
+Within this paper, we concentrate on the impacts of the LV CPT-even additive term $\frac{1}{2}\bar{\psi}H^{\mu\nu}\sigma_{\mu\nu}\psi$ in the Lagrangian, for which we will find the EH effective action.
 This is the aim we pursue in this paper.
@@
 $M$ $=$ $m+a_{\mu}\gamma^{\mu}+b_{\mu}\gamma^{\mu}\gamma_{5}+{\textstyle\frac{1}{2}}H^{\mu\nu}\sigma_{\mu\nu}.$ (2)
-Our aim consists in studying the one-loop low-energy effective action in the gauge sector, i.e.
-the LV EH action.
-Effectively we must obtain the contribution to the effective action of the lower order in a corresponding LV parameter, but including all orders in fields, from the following functional determinant
+Our aim consists in studying the one-loop low-energy effective action in the gauge sector, i.e., the LV EH action.
+Effectively, we must obtain the contribution to the effective action of the lower order in a corresponding LV parameter, but including all orders in fields, from the following functional determinant
 $\Gamma^{(1)}$ $=$ $-i{\rm Tr}\ln(i\Gamma^{\nu}D_{\nu}-M).$ (3)
-In our earlier studies EH1 , we have obtained the complete one-loop low-energy effective action (LEEA), that is, the EH action, for the lower orders in $b_{\mu}$ and $c_{\mu\nu}$, while the impact of $a^{\mu}$ is trivial in the Abelian case since $a^{\mu}$ is ruled out by the gauge transformation.
+Here and further, $\rm Tr$} is for the functional trace, while $\rm tr$} is for the simple matrix trace.
+In our earlier studies EH1 , we have obtained the complete one-loop low-energy effective action (LEEA), that is, the EH action, for the lower orders in $b_{\mu}$ and $c_{\mu\nu}$, while the impact of $a^{\mu}$ is trivial since the $a^{\mu}$ can be removed through a redefinition of a spinor field, see KosPic .
 As argued, and as it is very natural, the lower order in $b_{\mu}$ for the EH action must be the second one (the first order will yield either CFJ term or higher-derivative (HD) terms like Myers-Pospelov or HD CFJ-like terms, see JK ; HDLV , which do not match the EH structure).
 It is clear also that the EH structure cannot be generated by the first orders in $e^{\mu}$, $f^{\mu}$, $g^{\mu\nu\lambda}$ since they have odd numbers of indices, and by the first order in $d^{\mu\nu}$ for parity reasons, since $d^{\mu\nu}$ is a pseudotensor, so that to yield an observer Lorentz scalar form the quantum correction must include even orders in this parameter.
-Therefore, it is natural to study the possible one-loop effective action involving the first order in $H_{\mu\nu}$ which is expected to yield the EH form.
+Therefore, it is natural to study the possible one-loop effective action involving the first order in $H_{\mu\nu}$, which is expected to yield the EH form.
 While it is clear that the possible EH-like contributions of even orders in $F_{\mu\nu}$, like ${\rm tr}(HF^{2n})$ (e.g.
@@
 So, let us study these contributions to the EH action.
-We start with the expression (3) which in our case takes the form
+We start with the expression (3), which in our case takes the form
 […]
@@
 where we can use the key identity derived earlier in Schwinger:1951nm ; Dittrich:1985yb ; McA :
-$e^{isD^{2}}\delta^{4}(x-x^{\prime})|_{x=x^{\prime}}=\langle x|e^{isD^{2}}|x^{\prime}\rangle|_{x=x^{\prime}}=-\frac{i}{16\pi^{2}s^{2}}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$ (9)
-Indeed, it is clear that the term proportional to $D^{\lambda}$, in (7), yields a zero contribution to the trace being of odd order in Dirac matrices. Therefore, we are left with
+$e^{isD^{2}}\delta^{4}(x-x^{\prime})|_{x=x^{\prime}}=\langle x|e^{isD^{2}}|x^{\prime}\rangle|_{x=x^{\prime}}=-\frac{i}{16\pi^{2}s^{2}}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right),$ (9)
+here and further $F$ within the determinant is the matrix whose elements are given by $F_{\mu\nu}$.
+Indeed, it is clear that the term proportional to $D^{\lambda}$, in (7), yields a zero contribution to the trace being of odd order in Dirac matrices.
+Therefore, we are left with
 $\Gamma^{(1)}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}H_{\kappa\lambda}\sigma^{\kappa\lambda}\ e^{{\textstyle\frac{ies}{2}}F_{\mu\nu}\sigma^{\mu\nu}}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$ (10)
-It remains to expand the exponential in power series in $F_{\mu\nu}$:
+It remains to expand the exponential in a power series in $F_{\mu\nu}$:
 $\Gamma^{(1)}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}H_{\kappa\lambda}\sigma^{\kappa\lambda}\sum\limits_{n=0}^{\infty}\frac{({\textstyle\frac{ies}{2}}F_{\mu\nu}\sigma^{\mu\nu})^{2n+1}}{(2n+1)!}{\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$ (11)
@@
 ${\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right)=1+\frac{e^{2}s^{2}}{12}F_{\mu\nu}F^{\mu\nu}+{\cal O}(F^{4}).$ (12)
-Thus, after evaluating the trace, we obtain
-$\Gamma_{EH}$ $=$ $\frac{im}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}$ (13)
+Thus, after evaluating the trace of a product of $\sigma^{\mu\nu}$ matrices in the term of third order in $F_{\mu\nu}$ from (11), we obtain
+$\Gamma_{EH}$ $=$ $\frac{im}{32\pi^{2}}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}\ e^{ism^{2}}$ (13)
 $\times\left[\frac{ie^{3}s^{3}}{24}8H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-\frac{ie^{3}s^{3}}{48}(48H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-64H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa})\right].$
@@
 $\Gamma_{EH}=-\frac{e^{3}}{48\pi^{2}m^{3}}\int d^{4}x\ (H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-2H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa}).$ (14)
-In the next section we check this result with use of the Feynman diagrams.
-Now, in order to study the convergence of the series in (11), we can use the identity
+In the next section, we check this result with use of the Feynman diagrams.
+Now, to study the convergence of the series in (11), we can use the identity
 $\{\sigma^{\alpha\beta},\sigma^{\gamma\delta}\}=g^{\alpha\gamma}g^{\beta\delta}-g^{\alpha\delta}g^{\beta\gamma}+i\epsilon^{\alpha\beta\gamma\delta}\gamma_{5},$ (15)
@@
 So, we obtain
-$\Gamma^{(1)}$ $=$ $\frac{im}{16\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}$ (21)
+$\Gamma^{(1)}$ $=$ $\frac{im}{16\pi^{2}}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}$ (21)
 $\times\{H_{\kappa\lambda}F^{\kappa\lambda}[(F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}+(F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu})^{n}]$
@@
 $\times\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
-This sum can be evaluated. We note that the result is finite (indeed, the sum begins with $s^{3}$) as it must be by dimensional reasons.
+This sum can be evaluated. We note that the result is finite (indeed, the sum begins with $s^{3}$) as it must be for dimensional reasons.
 To perform the sum, we introduce the notation: $\Sigma=F_{\mu\nu}F^{\mu\nu}+iF_{\mu\nu}\tilde{F}^{\mu\nu}$, $\bar{\Sigma}=F_{\mu\nu}F^{\mu\nu}-iF_{\mu\nu}\tilde{F}^{\mu\nu}$. So, we can write
-$\Gamma^{(1)}$ $=$ $\frac{im}{16\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}[H_{\kappa\lambda}F^{\kappa\lambda}(\Sigma^{n}+\bar{\Sigma}^{n})-i\epsilon^{\kappa\lambda\rho\tau}H_{\kappa\lambda}F_{\rho\tau}(\Sigma^{n}-\bar{\Sigma}^{n})]$ (22)
+$\Gamma^{(1)}$ $=$ $\frac{im}{16\pi^{2}}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s^{2}}e^{ism^{2}}\sum\limits_{n=1}^{\infty}\frac{({\textstyle\frac{ies}{2}})^{2n+1}}{(2n+1)!}[H_{\kappa\lambda}F^{\kappa\lambda}(\Sigma^{n}+\bar{\Sigma}^{n})-i\epsilon^{\kappa\lambda\rho\tau}H_{\kappa\lambda}F_{\rho\tau}(\Sigma^{n}-\bar{\Sigma}^{n})]$ (22)
 $\times\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
 Using the Taylor series $\sum\limits_{n=1}^{\infty}\frac{A^{n}}{(2n+1)!}=\frac{\sinh\sqrt{A}-\sqrt{A}}{\sqrt{A}}$, we finally obtain
-$\Gamma^{(1)}$ $=$ $-\frac{em}{32\pi^{2}}{\rm tr}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s}e^{ism^{2}}\left[H_{\kappa\lambda}F^{\kappa\lambda}\left(\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\Sigma})-{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}{{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}+\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}})-{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}{{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}\right)\right$.
-(23)
+$\Gamma^{(1)}$ $=$ $-\frac{em}{32\pi^{2}}\int d^{4}x\int_{0}^{\infty}\frac{ds}{s}e^{ism^{2}}\left[H_{\kappa\lambda}F^{\kappa\lambda}\left(\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\Sigma})-{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}{{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}+\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}})-{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}{{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}\right)\right$.
 $\left.+i\epsilon^{\kappa\lambda\rho\tau}H_{\kappa\lambda}F_{\rho\tau}\left(\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\Sigma})-{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}{{\textstyle\frac{ies}{2}}\sqrt{\Sigma}}-\frac{\sinh({\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}})-{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}{{\textstyle\frac{ies}{2}}\sqrt{\bar{\Sigma}}}\right)\right]\ {\rm det}^{1/2}\left(\frac{esF}{\sinh(esF)}\right).$
 This is our final result.
-As we already noted, it is real, UV finite and begins with the cubic order in the stress tensor $F_{\mu\nu}$, which is nontrivial since earlier, in EH1 ; scalpt , only even orders in the stress tensor were shown to yield nontrivial contributions to the effective action.
+As we already noted, it is real; moreover, actually, (II) is nothing more as a compact presentation of (22), which involves only integer degrees of $\Sigma$ and $\bar{\Sigma}$ (hence there is no ambiguities caused by a multi-valued nature of a square root of a complex number in this result), so, the (II) must be interpreted as a power series in $\Sigma$ and $\bar{\Sigma}$.
+We note also that this result is UV finite and begins with the cubic order in the stress tensor $F_{\mu\nu}$, which is nontrivial since earlier, in EH1 ; scalpt , only even orders in the stress tensor were shown to yield nontrivial contributions to the effective action.
 III Three-point function
 In this section, we analyze the three-point function to calculate the one-loop EH effective action of the parameter $H_{\mu\nu}$ using the Feynman diagram approach.
-Our aim is to explicitly demonstrate that the result obtained reproduces the same structure that can be derived directly from the EH effective action (11).
+We aim to explicitly demonstrate that the result obtained reproduces the same structure that can be derived directly from the EH effective action (11).
 For this, given the Lagrangian
@@
 with
-$G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=2T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+2T_{H}^{\mu_{2}\mu_{1}\mu_{3}}(k_{2},k_{1},k_{3})$ (27)
-where $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ is subdivided into three graphs as follows:
+$G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=2T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+2T_{H}^{\mu_{2}\mu_{1}\mu_{3}}(k_{2},k_{1},k_{3}),$ (27)
+where, to first order in $H_{\mu\nu}$, both $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ and $T_{H}^{\mu_{2}\mu_{1}\mu_{3}}(k_{2},k_{1},k_{3})$ can be decomposed into three diagrams.
+In particular, for $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$, we have
 $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (28)
@@
 where $p_{1}^{\mu}=p^{\mu}-k_{1}^{\mu}$ and $p_{12}^{\mu}=p^{\mu}-k_{1}^{\mu}-k_{2}^{\mu}$.
-It is easy to see that $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ can be also obtained when we perform the ciclic interchanges:
+The corresponding Feynman diagrams for expressions (27) and (28) are shown in Figs.
+1 and 2, respectively, in the Appendix.
+It is easy to see that $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ can be obtained from $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$ by performing the cyclic interchanges:
 $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $T_{H1}^{\mu_{2}\mu_{3}\mu_{1}}(k_{2},k_{3},k_{1}),$ (30a)
 $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $T_{H1}^{\mu_{3}\mu_{1}\mu_{2}}(k_{3},k_{1},k_{2}).$ (30b)
-Therefore, we must focus our attention on the graph $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$, in which, by considering first the Feynman parameterization, we obtain
+Therefore, we must focus our attention on the graph $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$, in which, by first considering the Feynman parameterization, we obtain
 $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\int_{0}^{1}dx_{1}\int_{0}^{1-x_{1}}dx_{2}\int\frac{d^{4}p}{(2\pi)^{4}}\frac{3ixe^{3}}{(p^{2}-M^{2})^{4}}$ (31)
@@
 $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=\sum_{i=1}^{4}T_{H1ggi}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}T_{H1gki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}T_{H1kki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (35)
+where the explicit form of all relevant terms is given in Eq.
+(40) (see the Appendix).
+In this expression, we use the following notation: $T^{\mu_{1}\mu_{2}\mu_{3}}_{H1ggi}$ denotes terms in which all external fields are contracted with Minkowski metrics; in $T^{\mu_{1}\mu_{2}\mu_{3}}_{H1gki}$, two of the external fields are contracted with Minkowski metrics and the remaining one with an external momentum; finally, in $T^{\mu_{1}\mu_{2}\mu_{3}}_{H1kki}$ only one external field is contracted with a Minkowski metric and the remaining two with external momenta.
+In all these expressions, $i=1,2,3,4$ labels the corresponding contributions.
+Analogous definitions are also employed in the next equation.
+Let us now briefly discuss the question of photon double splitting in the collinear limit.
+If the incident on-shell photon has energy $E_{1}$ and momentum $\vec{k}_{1}$, then energy–momentum conservation requires that all momenta $\vec{k}_{i}$ must be aligned.
+Thus, the initial and final photons propagate collinearly, with orthogonal four-momenta $k_{i}^{\mu}k_{j\mu}=0$ (for more details, see Kostelecky:2002ue ).
+From the transversality condition, one finds $\epsilon_{i\mu}k_{j}^{\mu}=0$ (or $A_{\mu}(k_{i})k_{j}^{\mu}=0$).
+Hence, the contribution (40a) of order $\frac{1}{m}$ can yield a nonzero amplitude, while all other terms of (40) of order $\frac{1}{m^{3}}$ vanish in this limit.
+However, when we consider the other contributions coming from $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ through the permutations (30), all the contributions of order $\frac{1}{m}$ cancel each other, and thus photon double splitting does not occur.
+This fact was only mentioned in Kostelecky:2002ue , but not explicitly shown.
+For the photon triple splitting initially discussed in Kostelecky:2002ue , and later in FuMa , the contributions of order $\frac{1}{m^{2}}$ do not cancel each other when all the insertions of the coefficient $c_{\mu\nu}$ in the propagator are considered.
+Thus, after the collinear limit is taken into account, a nonzero amplitude is obtained, in contrast to what we observe here for the coefficient $H_{\mu\nu}$.
+The next step is to analyze the generation of the Lorentz-violating EH action for the first-order correction.
+For this, we must calculate $G_{H}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
+(27)), by initially calculating $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
+(28)), when we interchange the uncontracted indices as well as the momentum indices for obtaining $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
+(30)) from $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
+(40)).
+Proceeding in this way, we obtain
+$G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=\sum_{i=1}^{3}G_{Hggi}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{3}G_{Hgki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}G_{Hkki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (36)
+where the relevant terms are given in Eq. (41) in the Appendix. We observe that the contributions of order $\frac{1}{m}$ vanish, as expected.
+Thus, considering these results (36), the effective action (26), to first order in $H_{\mu\nu}$, takes the form
+$\Gamma_{EH}$ $=$ $\int d^{4}x\int d^{4}k_{1}d^{4}k_{2}d^{4}k_{3}\ e^{-i(k_{1}+k_{2}+k_{3})\cdot x}G_{H}(k_{1},k_{2},k_{3}),$ (37)
 where
+$G_{H}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}}{144\pi^{2}m^{3}}(H_{\kappa\lambda}F^{\kappa\lambda}(k_{1})F_{\mu\nu}(k_{2})F^{\mu\nu}(k_{3})+H_{\kappa\lambda}F^{\kappa\lambda}(k_{2})F_{\mu\nu}(k_{1})F^{\mu\nu}(k_{3})$ (38)
+$+H_{\kappa\lambda}F^{\kappa\lambda}(k_{3})F_{\mu\nu}(k_{1})F^{\mu\nu}(k_{2})-2H_{\kappa\lambda}F^{\lambda\mu}(k_{1})F_{\mu\nu}(k_{2})F^{\nu\kappa}(k_{3})$
+$-2F_{\kappa\lambda}(k_{1})H^{\lambda\mu}F_{\mu\nu}(k_{2})F^{\nu\kappa}(k_{3})-2F_{\kappa\lambda}(k_{1})F^{\lambda\mu}(k_{2})H_{\mu\nu}F^{\nu\kappa}(k_{3})),$
+with $F^{\mu\nu}(k_{1})=k_{1}^{\mu}A^{\nu}(k_{1})-k_{1}^{\nu}A^{\mu}(k_{1})$, and so on. Then, inverting the Fourier transform in Eq. (37), the Lorentz-violating Euler-Heisenberg action becomes
+$\Gamma_{EH}=-\frac{e^{3}}{48\pi^{2}m^{3}}\int d^{4}x\ (H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-2H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa}),$ (39)
+where now $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$.
+We note that, using this definition of $F_{\mu\nu}$, this result can be easily rewritten in the form (26), that is, in terms of the vector potential $A_{\mu}$.
+This result exactly coincides with the expression (14) obtained above with the use of the proper time method, which confirms the validity of our calculations.
+IV Summary
+We calculated the one-loop EH effective action for the LV spinor QED characterized by the LV parameter $H_{\mu\nu}$ of dimension one, which ensures finiteness of our result.
+Its unusual feature consists in the fact that, unlike the standard EH effective action (including many cases with the presence of LV terms, see EH1 ; scalpt ), our one-loop expression involves only odd orders in $F_{\mu\nu}$.
+Another important observation regarding our result is that it involves the first order in the LV parameter, while in many other studies of the EH effective action, the result is of the second order in corresponding LV parameters (see e.g.
+EH1 ; scalpt ).
+Therefore, our result may be the dominant contribution due to the well-known smallness of LV parameters datatables .
+We explicitly demonstrated that the third-order result obtained through the EH framework coincides with that calculated through the Feynman diagrams approach.
+Effectively, we succeeded in obtaining results analogous to those found in FuMa ; Kostelecky:2002ue for our theory.
+We plan to study phenomenological implications of our results in further studies.
+Another interesting problem consists in obtaining the EH effective action depending on other minimal LV parameters presented in the model (1), namely, $d^{\mu\nu},e^{\mu},f^{\mu},g^{\mu\nu\lambda}$.
+We note that these calculations will be more complicated than those performed here and in EH1 , since in the presence of these parameters, calculations of more involved traces will be necessary.
+We plan to perform these studies in our forthcoming papers.
+Acknowledgments.
+The work of T.
+M.
+has been partially supported by the CNPq project No.
+309360/2025-0 and FAPEAL project No.
+E:60030.0000002341/2022.
+The work of A.
+Yu.
+P.
+has been partially supported by the CNPq project No.
+303777/2023-0.
+Appendix
+The relevant contributions to $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ explicitly look like
 $T_{H1gg1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $\frac{ie^{3}H_{\alpha\beta}}{48\pi^{2}m}(g^{\mu_{1}\mu_{2}}((k_{2}^{\beta}-2k_{1}^{\beta})g^{\alpha\mu_{3}}+(2k_{1}^{\alpha}-k_{2}^{\alpha})g^{\beta\mu_{3}})+g^{\mu_{1}\mu_{3}}(k_{2}^{\alpha}g^{\beta\mu_{2}}-k_{2}^{\beta}g^{\alpha\mu_{2}})$ (40a)
@@
 $-(5k_{1}^{\beta}+2k_{2}^{\beta})g^{\alpha\mu_{2}})+3k_{2}^{\mu_{2}}(k_{1}^{\beta}g^{\alpha\mu_{1}}-k_{1}^{\alpha}g^{\beta\mu_{1}}))).$
-Let us now briefly discuss the question of photon double splitting in the collinear limit.
-If the incident on-shell photon has energy $E_{1}$ and momentum $\vec{k}_{1}$, then energy–momentum conservation requires that all momenta $\vec{k}_{i}$ must be aligned.
-Thus, the initial and final photons propagate collinearly, with orthogonal four-momenta $k_{i}^{\mu}k_{j\mu}=0$ (for more details, see Kostelecky:2002ue ).
-From the transversality condition, one finds $\epsilon_{i\mu}k_{j}^{\mu}=0$ (or $A_{\mu}(k_{i})k_{j}^{\mu}=0$).
-Hence, the contribution (36a) of order $\frac{1}{m}$ can yield a nonzero amplitude, while all other terms of (36) of order $\frac{1}{m^{3}}$ vanish in this limit.
-However, when we consider the other contributions coming from $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ through the permutations (30), all the contributions of order $\frac{1}{m}$ cancel each other, and thus photon double splitting does not occur.
-This fact was only mentioned in Kostelecky:2002ue , but not explicitly shown.
-For the photon triple splitting initially discussed in Kostelecky:2002ue , and later in FuMa , the contributions of order $\frac{1}{m^{2}}$ do not cancel each other when all the insertions of the coefficient $c_{\mu\nu}$ in the propagator are considered.
-Thus, after the collinear limit is taken into account, a nonzero amplitude is obtained, in contrast to what we observe here for the coefficient $H_{\mu\nu}$.
-The next step is to analyze the generation of the Lorentz-violating EH action, for the first order correction.
-For this, we must calculate $G_{H}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
-(27)), by initially calculating $T_{H}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
-(28)), when we interchange the uncontracted indices as well as the momentum indices for obtaining $T_{H2}^{\mu_{1}\mu_{2}\mu_{3}}$ and $T_{H3}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
-(30)) from $T_{H1}^{\mu_{1}\mu_{2}\mu_{3}}$ (Eq.
-(36)).
-Proceeding in this way, we obtain
-$G_{H}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})=\sum_{i=1}^{3}G_{Hggi}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{3}G_{Hgki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})+\sum_{i=1}^{4}G_{Hkki}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3}),$ (37)
-where
+Then, the contributions to the effective action are given by
 $G_{Hgg1}^{\mu_{1}\mu_{2}\mu_{3}}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}H_{\alpha\beta}}{12\pi^{2}m^{3}}g^{\mu_{1}\mu_{2}}(k_{1}^{\alpha}(k_{2}^{2}-k_{1}\cdot k_{2})g^{\beta\mu_{3}}+k_{2}^{\alpha}(k_{1}^{2}-k_{1}\cdot k_{2})g^{\beta\mu_{3}}+g^{\alpha\mu_{3}}$ (41a)
@@
 $+2k_{2}^{\mu_{2}}(k_{1}^{\beta}g^{\alpha\mu_{1}}-k_{1}^{\alpha}g^{\beta\mu_{1}}))).$
-We observe that the contributions of order $\frac{1}{m}$ vanish, as expected.
-Thus, by considering these results (37), the effective action (26) takes the form
-$\Gamma_{EH}$ $=$ $\int d^{4}x\int d^{4}k_{1}d^{4}k_{2}d^{4}k_{3}\ e^{-i(k_{1}+k_{2}+k_{3})\cdot x}G_{H}(k_{1},k_{2},k_{3}),$ (39)
-where
-$G_{H}(k_{1},k_{2},k_{3})$ $=$ $-\frac{ie^{3}}{144\pi^{2}m^{3}}(H_{\kappa\lambda}F^{\kappa\lambda}(k_{1})F_{\mu\nu}(k_{2})F^{\mu\nu}(k_{3})+H_{\kappa\lambda}F^{\kappa\lambda}(k_{2})F_{\mu\nu}(k_{1})F^{\mu\nu}(k_{3})$ (40)
-$+H_{\kappa\lambda}F^{\kappa\lambda}(k_{3})F_{\mu\nu}(k_{1})F^{\mu\nu}(k_{2})-2H_{\kappa\lambda}F^{\lambda\mu}(k_{1})F_{\mu\nu}(k_{2})F^{\nu\kappa}(k_{3})$
-$-2F_{\kappa\lambda}(k_{1})H^{\lambda\mu}F_{\mu\nu}(k_{2})F^{\nu\kappa}(k_{3})-2F_{\kappa\lambda}(k_{1})F^{\lambda\mu}(k_{2})H_{\mu\nu}F^{\nu\kappa}(k_{3})),$
-with $F^{\mu\nu}(k_{1})=k_{1}^{\mu}A^{\nu}(k_{1})-k_{1}^{\nu}A^{\mu}(k_{1})$, and so on. Then, inverting the Fourier transform in Eq. (39), the Lorentz-violating Euler-Heisenberg action becomes
-$\Gamma_{EH}=-\frac{e^{3}}{48\pi^{2}m^{3}}\int d^{4}x\ (H_{\kappa\lambda}F^{\kappa\lambda}F_{\mu\nu}F^{\mu\nu}-2H_{\kappa\lambda}F^{\lambda\mu}F_{\mu\nu}F^{\nu\kappa}),$ (41)
-where now $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$.
-This result exactly coincides with the expression (14) obtained above with use of the proper time methods which confirms the validity of our calculations.
-IV Summary
-We calculated the one-loop EH effective action for the LV spinor QED characterized by the LV parameter $H_{\mu\nu}$ of dimension one, which ensures finiteness of our result.
-Its unusual feature consists in the fact that, unlike the standard EH effective action (including many cases with the presence of LV terms, see EH1 ; scalpt ), our one-loop expression involves only odd orders in $F_{\mu\nu}$.
-Another important observation regarding our result is that it involves the first order in the LV parameter while in many other studies of the EH effective action the result is of the second order in corresponding LV parameters (see e.g.
-EH1 ; scalpt ).
-Therefore, our result may be the dominant contribution due to the well-known smallness of LV parameters datatables .
-We explicitly demonstrated that the third-order result obtained through the EH framework coincides with that calculated through the Feynman diagrams approach.
-Effectively, we succeeded in obtaining results analogous to those found in FuMa ; Kostelecky:2002ue for our theory.
-We plan to study phenomenological impacts of our results in further studies.
-Another interesting problem consists in obtaining the EH effective action depending on other minimal LV parameters presented in the model (1), namely, $d^{\mu\nu},e^{\mu},f^{\mu},g^{\mu\nu\lambda}$.
-We note that these calculations will be more complicated than those performed here and in EH1 , since in the presence of these parameters, calculations of more involved traces will be necessary.
-We plan to perform these studies in our forthcoming papers.
-Acknowledgments.
-The work of T.
-M.
-has been partially supported by the CNPq project No.
-309360/2025-0 and FAPEAL project No.
-E:60030.0000002341/2022.
-The work of A.
-Yu.
-P.
-has been partially supported by the CNPq project No.
-303777/2023-0.
+The corresponding Feynman diagrams are given below.
+Figura 1: The two contributions to the three-point function.
+Figura 2: The first contribution with insertions of $H_{\mu\nu}$ in the propagator.
 Referências